# Chemistry-Inspired Diffusion with Non-Diffe-rentiable Guidance

### v1

@@ 5 Analyses
In Appendix F, we discuss acceleration methods for ChemGuide (Appendix F.1); generalization analyses of property regressor (Appendix F.2) and conditional diffusion model (Appendix F.3); the effect of guidance scales for noisy/clean guidance (Appendix F.4); the effect of optimization steps for clean guidance (Appendix  F.5).
@@ 6 Related Work
Due to space limits, we present the discussion of some related works in Appendix B.
@@ 7 Conclusion & Future Work
Appendix A Implementation Details
@@
  • energy above ground state (Eh): it is the difference between the energy before and after using xTB or DFT (specified in the context) to optimize the geometry of the molecules, and a smaller energy above ground state indicates the geometry is closer to the stable ground state of the molecule; it is the lower the better.
@@
We use one step of guidance/optimization for our experiments (e.g., clean guidance in Section 3.1) by default. Our implementation is available at https://anonymous.4open.science/r/ChemGuide.
Appendix B Related Work
@@
Previous works focus on generating molecules as SMILES strings (Kusner et al., 2017; Gómez-Bombarelli et al., 2018; Segler et al., 2018), and 2D graphs (Li et al., 2018; Jo et al., 2022), where various models, such as VAE (Simonovsky & Komodakis, 2018; Jin et al., 2018; 2019), GAN (De Cao & Kipf, 2018; Prykhodko et al., 2019), diffusion models (Vignac et al., 2022; Kong et al., 2023; Zhao et al., 2024), have been proposed to serve as the generative backbone. To model the spatial structure of molecules, G-Schnet (Gebauer et al., 2019) and G-SphereNet (Luo & Ji, 2022) use autoregressive models to generate the 3D coordinates of the molecules, however, they are less powerful compared with diffusion models. Hoogeboom et al. (2022) propose to use the equivariant diffusion model (EDM) for 3D molecule generation, which utilizes an equivariant graph neural network (Satorras et al., 2021) to model the molecules as a 3D graph, with coordinates and atom types as its node features. GeoLDM (Xu et al., 2023) further extends EDM to latent diffusion, and has shown improvements in stability and validity of the generated molecules. However, to achieve conditional generation, both EDM and GeoLDM need to be re-trained with labels, where the target property value is appended to the feature space as auxiliary information, which guides the generation toward fulfilling certain molecular property requirements.
@@
Guided diffusion has shown promising results in vision (Bansal et al., 2023), where the guidance can be provided by texts (Rombach et al., 2022), poses and edges (Zhang et al., 2023), classifiers (Dhariwal & Nichol, 2021), etc. Similar approaches are adapted for diffusion-based molecule generation, where explicitly guided diffusion (Hoogeboom et al., 2022; Xu et al., 2023) trains a conditional model with labels, and implicitly guided diffusion (Vignac et al., 2022; Bao et al., 2022) combines an unconditional model with a conditional property network, which provides guidance as gradients to the diffusion process at inference. Gruver et al. (2023) propose guided discrete diffusion for multi-objective protein sequence optimization, and Han et al. (2024) achieve multi-property guidance by modeling property relations as probabilistic graphs. Weiss et al. (2023) investigate guided diffusion and demonstrate the capability to design molecules beyond the seen property distribution, while MOOD (Lee et al., 2023) and GCD (Klarner et al., 2024) incorporate out-of-distribution control to improve generalization of the property predictor. However, unlike ChemGuide, all of the above works require a labeled dataset to train a property network; thus, inevitably constrained by the generalization ability of the neural networks.
Appendix C GFN2-xTB
@@
Its accuracy and efficiency come strictly from the element-specific and global parameters for all elements up to radon (Z = 86) Bannwarth et al. (2019), hence the semi-empiricism. The pre-computed tight-binding parameters and empirical corrections are utilized to approximate the electronic structure and calculate energy contributions efficiently.
Appendix D Motivation for Stability Optimization
@@
Appendix E Chemistry reasoning behind the bilevel property optimization
@@
Appendix F Analyses
F.1 Acceleration Methods for ChemGuide
@@
F.2 Generalization Analysis of Neural Regressors

### v2

@@ 5 Analyses
In Appendix H, we discuss acceleration methods for ChemGuide (Section H.1); different guidance steps (Section H.2); neural guidance for stability optimization (Section H.3); generalization analyses of property regressor (Section H.4) and conditional diffusion model (Section H.5); the effect of guidance scales for noisy/clean guidance (Section H.6); the effect of optimization steps for clean guidance (Section H.7).
@@ 6 Related Work
Due to space limits, we present the discussion of related works in Appendix C.
@@ 7 Conclusion & Future Work
8 Acknowledgments
We are grateful to our anonymous ICLR reviewers, with whose help this work has been greatly improved. The work was partially supported by the NSF Award 2307698.
Appendix A List of Contents
We detail the contents in the appendix below.
  • Implementation Details (Section B) includes the hardware and running time, the models, the evaluation metrics for our experiments, and the link to the code repository.
  • Related Work (Section C) includes related literature on molecule generation and guided diffusion.
  • GFN2-xTB (Section D) introduces the method used by ChemGuide to perform quantum chemistry calculation.
  • Motivation for Stability Optimization (Section E) explains the importance and motivation of stability optimization for molecule generation.
  • Motivation for Gradient Estimation with SPSA (Section F) explains the motivation of choosing SPSA as the zeroth-order optimization method to estimate gradients.
  • Chemistry Reasoning behind Bilevel Property Optimization (Section G) explains the chemistry intuition for the bilevel optimization framework on molecular property and stability.
  • Analyses (Section H) discusses acceleration methods for ChemGuide (Section H.1); different guidance steps (Section H.2); neural guidance for stability optimization (Section H.3); generalization analyses of property regressor (Section H.4) and conditional diffusion model (Section H.5); the effect of guidance scales for noisy/clean guidance (Section H.6); the effect of optimization steps for clean guidance (Section H.7).
  • Other Optimization Algorithms (Section I) discusses the possibility of conditional generation and stability optimization with evolutionary algorithm, and how it can be used together with ChemGuide.
  • Bilevel Optimization with Clean Guidance (Section J) discusses using clean guidance and ChemGuide in the bilevel optimization framework.
  • Comprehensive Results on Non-differentiable Oracle Guidance (Section K) presents detailed results of ChemGuide for stability and molecular property optimization.
  • Comprehensive Results on Neural Guidance (Section L) presents detailed results of noisy and clean guidance for molecular property optimization.
  • Comprehensive Results on Bilevel Optimization (Section M) presents detailed results for molecular property and stability optimization.
  • Molecule Visualization (Section N) visualizes the generated molecules.
Appendix B Implementation Details
@@
  • energy above ground state (Eh): it is the difference between the energy before and after using xTB or DFT (specified in the context) to optimize the geometry of the generated molecules, and a smaller energy above ground state indicates the geometry is closer to the stable ground state of the molecule; it is the lower the better.
@@
We use one step of guidance/optimization for our experiments (e.g., clean guidance in Section 2) by default. Our implementation is available at https://github.com/A-Chicharito-S/ChemGuide.
Appendix C Related Work
@@
Previous works focus on generating molecules as SMILES strings (Kusner et al., 2017; Gómez-Bombarelli et al., 2018; Segler et al., 2018), and 2D graphs (Li et al., 2018; Jo et al., 2022), where various models, such as VAE (Simonovsky & Komodakis, 2018; Jin et al., 2018; 2019), GAN (De Cao & Kipf, 2018; Prykhodko et al., 2019), diffusion models (Vignac et al., 2022; Kong et al., 2023; Zhao et al., 2024a), have been proposed to serve as the generative backbone. To model the spatial structure of molecules, G-Schnet (Gebauer et al., 2019) and G-SphereNet (Luo & Ji, 2022) use autoregressive models to generate the 3D coordinates of the molecules, however, they are less powerful compared with diffusion models. Hoogeboom et al. (2022) propose to use the equivariant diffusion model (EDM) for 3D molecule generation, which utilizes an equivariant graph neural network (Satorras et al., 2021) to model the molecules as a 3D graph, with coordinates and atom types as its node features. GeoLDM (Xu et al., 2023) further extends EDM to latent diffusion, and has shown improvements in stability and validity of the generated molecules. However, to achieve conditional generation, both EDM and GeoLDM need to be re-trained with labels, where the target property value is appended to the feature space as auxiliary information, which guides the generation toward fulfilling certain molecular property requirements.
@@
Guided diffusion has shown promising results in vision (Bansal et al., 2023), where the guidance can be provided by texts (Rombach et al., 2022), poses and edges (Zhang et al., 2023), classifiers (Dhariwal & Nichol, 2021), etc. Similar approaches are adapted for diffusion-based molecule generation, where explicitly guided diffusion (Hoogeboom et al., 2022; Xu et al., 2023) trains a conditional model with labels, and implicitly guided diffusion (Vignac et al., 2022; Bao et al., 2022) combines an unconditional model with a conditional property network, which provides guidance as gradients to the diffusion process at inference. Gruver et al. (2023) propose guided discrete diffusion for multi-objective protein sequence optimization, and Han et al. (2024) achieve multi-property guidance by modeling property relations as probabilistic graphs. Weiss et al. (2023) investigate guided diffusion and demonstrate the capability to design molecules beyond the seen property distribution, while MOOD (Lee et al., 2023) and GCD (Klarner et al., 2024) incorporate out-of-distribution control to improve generalization of the property predictor. However, unlike ChemGuide, all of the above works require a labeled dataset to train a property network; thus, inevitably constrained by the generalization ability of the neural networks. Beyond molecule optimization, diffusion guidance has been discussed in the broader literature, such as protein design (Wu et al., 2023), large language models (Zhao et al., 2024b), and music generation (Huang et al., 2024).
Appendix D GFN2-xTB
@@
Its accuracy and efficiency come strictly from the element-specific and global parameters for all elements up to radon (Z = 86) (Bannwarth et al., 2019), hence the semi-empiricism. The pre-computed tight-binding parameters and empirical corrections are utilized to approximate the electronic structure and calculate energy contributions efficiently.
Appendix E Motivation for Stability Optimization
@@
Appendix F Motivation for Gradient Estimation with SPSA
To achieve equivariance to transformations (Eq. 4), the 3D diffusion models must satisfy the zero center gravity requirement that $\mathbf{z}_{\mathbf{x},t}\in\mathbb{R}^{N\times 3}$ should have zero-mean over $N$ atoms on the $x,y,z$ axis. To derive guidance, one can estimate $\nabla_{{\bm{z}}_{t}}\mathcal{F}({\bm{z}}_{t})$ in the gradient using finite-different approximation (from Eq. 14) as:
$\nabla_{{\bm{z}}_{t}}\mathcal{F}({\bm{z}}_{t})$ $\approx\lim_{\zeta\rightarrow 0}\frac{\mathcal{F}([{\bm{z}}_{\mathbf{x},t}+\zeta\mathbf{1}_{N\times 3},{\bm{z}}_{\mathbf{h},t}])-\mathcal{F}([{\bm{z}}_{\mathbf{x},t}-\zeta\mathbf{1}_{N\times 3},{\bm{z}}_{\mathbf{h},t}])}{2\zeta}$ (19)
which however violates zero center gravity as the perturbed variable ${\bm{z}}_{\mathbf{x},t}\pm\zeta\mathbf{1}_{N\times 3}$ is shifted by ±$\pm\zeta\mathbf{1}_{N\times 3}$ over the zero-mean of ${\bm{z}}_{\mathbf{x},t}$.
To obey the zero-mean requirements, one can employ SPSA (Simultaneous Perturbation Stochastic Approximation) (Spall, 1992) by perturbing ${\bm{z}}_{\mathbf{x},t}$ with a zero-mean random variable $\Delta_{t}$ as follows.
$\nabla_{{\bm{z}}_{t}}\mathcal{F}({\bm{z}}_{t})$ $\approx\lim_{\zeta\rightarrow 0}\frac{\mathcal{F}([{\bm{z}}_{\mathbf{x},t}+\zeta\Delta_{t},{\bm{z}}_{\mathbf{h},t}])-\mathcal{F}([{\bm{z}}_{\mathbf{x},t}-\zeta\Delta_{t},{\bm{z}}_{\mathbf{h},t}])}{2\zeta\Delta_{t}}$ (20)
where we abuse the notation by using dividing $\Delta_{t}$ to denote the element-wise operation. Spall (1992) assumes ${\bm{z}}_{\mathbf{x},t}$ to have bounded inverse moments, which excludes $\Delta_{t}$ being Gaussian. However, the Gaussian perturbation $[{\bm{z}}_{\mathbf{x},t}\pm\zeta\mathbf{U},{\bm{z}}_{\mathbf{h},t}]$ with $\mathbf{U}\in\mathbb{R}^{N\times 3}\sim\mathcal{N}(0,1)$ would be favored, since it is one of the commonly used distributions that obey zero central gravity with $\mathbb{E}[\zeta\mathbf{U}]=0$, while only introduce small computational overhead compared to finite-different methods, as sampling from $\mathcal{N}(0,1)$ and matrix multiplication are fast in modern machine learning practices on GPUs.
To achieve so, one can adopt the random gradient-free oracles (Nesterov & Spokoiny, 2017) to estimate the gradient for differentiable $f:E\rightarrow R$. Given the random vector $u\in E$ being Gaussian with correlation operator $B^{-1}$ (e.g., $B=I$) and $\mu\geq 0$, the gradient at $x\in E$ is estimated as follows.
$\hat{g}_{u}(x)=\frac{f(x+\mu u)-f(x-\mu u)}{2\mu}\cdot Bu$ (21)
which leads to our estimation for the guidance gradient in Eq. 15.
Appendix G Chemistry Reasoning behind Bilevel Property Optimization
@@
Appendix H Analyses
H.1 Acceleration Methods for ChemGuide
@@
H.2 Effect of Guidance Steps
As mentioned in Appendix B, adding guidance to the last 400 of the 1000 diffusion steps is time-consuming, so we explore relaxations in this section, where we add guidance to later steps such as the last 300, 200, and 100 steps. The results are presented in Table 8 for QM9 and Table 9 for GEOM. It is observed that more guidance steps yield better results in terms of force RMS and energy above ground state while the effect on validity and stability does not seem to be related to the number of guidance steps. This is reasonable because we are optimizing the force during the diffusion process, so adding more guidance steps provides more guidance towards lower force and smaller energy above ground state.
Table 8: Metrics of 500 generated molecules from QM9 using GeoLDM with non-differentiable oracle (i.e. xTB) guidance with various guidance steps (100, 200, 300, 400) and guidance scales (0.0001, 0.001, 0.01, 0.1, 1.0). * and bold denote the overall best result and our best result, respectively. Percentage changes between our results and GeoLDM are shown in parentheses.
Metric Guidance Scale (100 Guidance Steps) Baseline
0.0001 0.001 0.01 0.1 1.0 EDM GeoLDM
Force RMS (Eh/Bohr) 0.0110 (-0.88$\downarrow$) 0.0110 (-0.88$\downarrow$) 0.0110 0.0113 0.0120 0.0114 0.0111
Validity 90.40% 90.40% 90.40% 90.80% (1.00$\uparrow$) 90.60% 86.60% 89.80%
Uniqueness 100.00% 100.00% 100.00% 100.00% 100.00% 100.00% 100.00%
Atom Stability 98.94% (0.01$\uparrow$) 98.94% (0.01$\uparrow$) 98.94% (0.01$\uparrow$) 98.91% 98.92% 98.53% 98.93%
Molecule Stability 89.40% 89.40% 89.40% 89.60% (0.80$\uparrow$) 89.60% (0.80$\uparrow$) 83.60% 88.80%
Energy above ground state (Eh) 0.0051 0.0051 (0.87$\uparrow$) 0.0051 0.0055 0.0057 0.0072 0.0050
Metric Guidance Scale (200 Guidance Steps) Baseline
0.0001 0.001 0.01 0.1 1.0 EDM GeoLDM
Force RMS 0.0119 0.0106 (-4.28$\downarrow$) 0.0119 0.0114 0.0116 0.0114 0.0111
Validity 90.00% 91.00% (1.20$\uparrow$) 90.00% 90.60% 90.60% 86.60% 89.80%
Uniqueness 100.00% 100.00% 100.00% 100.00% 100.00% 100.00% 100.00%
Atom Stability 98.87% 98.72% 98.87% 98.90% (-0.03$\downarrow$) 98.86% 89.53% 98.93%
Molecule Stability 89.00% 88.40% 89.00% 89.40% (0.60$\uparrow$) 89.00% 83.60% 88.80%
Energy above ground state (Eh) 0.0054 0.0045 (-9.32$\downarrow$) 0.0054 0.0049 0.0050 0.0072 0.0050
Metric Guidance Scale (300 Guidance Steps) Baseline
0.0001 0.001 0.01 0.1 1.0 EDM GeoLDM
Force RMS 0.0109 0.0109 0.0107 (-3.98$\downarrow$) 0.0108 0.0115 0.0114 0.0111
Validity 92.80% (3.00$\uparrow$) 92.80% (3.00$\uparrow$) 92.60% 92.20% 88.80% 86.60% 89.80%
Uniqueness 99.57% 100.00% 99.58% 99.36% 100.00% 100.00% 100.00%
Atom Stability 98.93% 99.09% (0.15$\uparrow$) 99.00% 98.99% 98.88% 98.53% 98.93%
Molecule Stability 89.80% 90.80% (2.00$\uparrow$) 89.60% 89.60% 88.00% 83.60% 88.80%
Energy above ground state (Eh) 0.0049 0.0052 0.0045 (-10.64$\downarrow$) 0.0046 0.0056 0.0072 0.0050
Metric Guidance Scale (400 Guidance Steps) Baseline
0.0001 0.001 0.01 0.1 1.0 EDM GeoLDM
Uniqueness 100.00% 100.00% 100.00% 100.00% 100.00% 100.00% 100.00%
Table 9: Metrics of 500 generated molecules from GEOM using GeoLDM with non-differentiable oracle (i.e. xTB) guidance with various guidance steps (100, 200, 300, 400) and guidance scales (0.0001, 0.001, 0.01, 0.1, 1.0). * and bold denote the overall best result and our best result, respectively. Percentage changes between our results and GeoLDM are shown in parentheses.
Metric Guidance Scale (100 Guidance Steps) Baseline
0.0001 0.001 0.01 0.1 1.0 EDM GeoLDM
Force RMS (Eh/Bohr) 0.0434 0.0434 0.0432 0.0430 0.0414 (-13.36$\downarrow$) 0.0742 0.0478
Validity 49.20% 49.40% 49.20% 50.00% (1.00$\uparrow$) 48.00% 46.40% 49.00%
Uniqueness 100.00% 100.00% 100.00% 100.00% 100.00% 100.00% 100.00%
Atom Stability 84.70% 84.70% (0.17$\uparrow$) 84.68% 84.67% 84.67% 81.22% 84.53%
Energy above Ground State (Eh) 0.2043 0.2039 0.2031 (-9.66$\downarrow$) 0.2049 0.2037 0.3742 0.2248
Metric Guidance Scale (200 Guidance Steps) Baseline
0.0001 0.001 0.01 0.1 1.0 EDM GeoLDM
Force RMS (Eh/Bohr) 0.0433 0.0425 (-11.13$\downarrow$) 0.0445 0.0433 0.0476 0.0742 0.0478
Validity 48.80% 48.40% 50.20% 47.80% 51.00% (2.00$\uparrow$) 46.40% 49.00%
Uniqueness 100.00% 100.00% 100.00% 100.00% 100.00% 100.00% 100.00%
Atom Stability 84.31% 84.28% 84.28% 84.45% (-0.08$\downarrow$) 83.91% 81.22% 84.53%
Energy above Ground State (Eh) 0.2224 0.2181 0.2254 0.2142 (-4.72$\downarrow$) 0.2297 0.3742 0.2248
Metric Guidance Scale (300 Guidance Steps) Baseline
0.0001 0.001 0.01 0.1 1.0 EDM GeoLDM
Force RMS (Eh/Bohr) 0.0462 0.0471 0.0468 0.0440 (-7.96$\downarrow$) 0.0478 0.0742 0.0478
Validity 47.20% 48.50% 48.60% 51.80% (2.80$\uparrow$) 50.00% 46.40% 49.00%
Uniqueness 100.00% 100.00% 100.00% 100.00% 100.00% 100.00% 100.00%
Atom Stability 83.73% 83.59% 83.75% 83.80% (-0.73$\downarrow$) 83.38% 81.22% 84.53%
Energy above Ground State (Eh) 0.2190 0.2248 0.2223 0.2071 (-7.88$\downarrow$) 0.2520 0.3742 0.2248
Metric Guidance Scale (400 Guidance Steps) Baseline
0.0001 0.001 0.01 0.1 1.0 EDM GeoLDM
Uniqueness 100.00% 100.00% 100.00% 100.00% 100.00% 100.00% 100.00%
H.3 ChemGuide vs. Neural Guidance on Force
In this section, we replace the non-differentiable oracle (i.e. xTB) guidance with a neural network that predicts energy and force. We employ a machine learning interatomic potential (MLIP), AIMNet2 (Anstine et al., 2024), a recent neural network model trained on 20 million hybrid quantum chemical calculations to efficiently predict energy, force, and properties. By combining ML-parameterized short-range and physics-based long-range terms, AIMNet2 achieves better generalization across diverse molecules. We sample 500 molecules from QM9 and 50 molecules from GEOM, and present the results in Table 10 for QM9 and Table 11 for GEOM, for reference and comparison, we also present the metrics of ChemGuide with the best scales.
Using AIMNet2 as guidance doesn’t improve over xTB in terms of time, because it introduces a prohibitive memory constraint that makes sampling almost linear. Recall from Section 2, the molecule position is represented as $[N,3]$; thus, the force is $[N,3]$, and the derivative (i.e. Hessian) of the force on the position is $[N,3,N,3]$, which is required when calculating the guidance gradient via backpropagation. It requires a large amount of GPU memory and is problematic when sampling large molecules (e.g., GEOM). In ChemGuide, only force is required to estimate gradient, and 100 molecules are sampled at a time on a 48GiB GPU, however, with AIMNet2 guidance only 5 molecules can be sampled at a time on the same machine. This is because calculating Hessian scales up the memory by $N\times 3$, which is a major drawback of using neural networks on force guidance as it requires more memory.
For QM9, we sample 500 molecules, and using AIMNet2 as guidance yields much worse results in all metrics than ChemGuide. For GEOM, we sample 50 molecules due to the memory constraint, and AIMNet2 achieves compatible results as ChemGuide for guidance, where the force RMS, validity, and stability are close.
However, on GEOM, ChemGuide achieves much lower energy above ground state compared to AIMNet2 guidance, indicating that ChemGuide generated molecules closer to their corresponding ground states. Recall that in the energy curve of a molecule, the x-axis and the y-value refer to the molecule configuration and its energy, where the derivative of energy gives the force, and we aim to optimize the molecule towards its ground state (i.e. the global minimum of the energy curve). Hence, although AIMNet2 guidance reduces the force similarly to ChemGuide, it is farther from the global energy minimum, which demonstrates that guidance directly from quantum chemistry provides more intrinsic optimization than neural networks. Additionally, we sample 500 molecules in Table 2 while we only sample 50 molecules for AIMNet, which might not be comparable as smaller populations are more likely to be biased.
Figure 6: Cosine similarity between the guidance gradients of AIMNet and ChemGuide of 10 molecules sampled on QM9 for the last 400 steps.
To better understand the difference, we plot the cosine similarities of the gradients from AIMNet2 and ChemGuide over 400 guidance steps on 10 molecules from QM9, as shown in Figure 6. The results show that the cosine similarities are small and concentrated around zero, suggesting that the gradients from AIMNet2 and ChemGuide are nearly orthogonal. This observation helps explain the significant performance differences reported in Table 10: ChemGuide achieves much better results than AIMNet2-guided optimization on QM9 because ChemGuide provides more effective and stable guidance, whereas the guidance from AIMNet2 negatively impacts stability.
Table 10: Metrics of 500 generated molecules from QM9 using GeoLDM with neural network guidance (i.e. AIMNet2) on force. * and bold denote the overall best result and our best result, respectively. Percentage changes between our results and GeoLDM are shown in parentheses. For ease of comparison, we report the results using the best scale of ChemGuide from Table 1.
Metric Guidance Scale ChemGuide Baseline
0.0001 0.001 0.01 0.1 1.0 scale = 0.0001 EDM GeoLDM
Force RMS (Eh/Bohr) 0.0113 0.0113 0.0112 0.0114 0.0110 (-1.00$\downarrow$) 0.0104 (-6.76$\downarrow$) 0.0114 0.0111
Validity 89.20% (-0.60$\downarrow$) 89.20% (-0.60$\downarrow$) 89.00% 88.60% 89.00% 91.40% (1.60$\uparrow$) 86.60% 89.80%
Uniqueness 100.00% 100.00% 100.00% 100.00% 100.00% 100.00% 100.00% 100.00%
Atom Stability 98.72% 98.72% 98.71% 98.74% 98.75% (-0.18$\downarrow$) 99.02% (0.09$\uparrow$) 98.53% 98.93%
Molecule Stability 87.80% 87.80% 87.60% 87.80% 88.00% (-0.80$\downarrow$) 90.60% (1.80$\uparrow$) 83.60% 88.80%
Energy above ground state (Eh) 0.0056 0.0057 0.0056 0.0059 0.0055 (9.30$\uparrow$) 0.0042 (-15.78$\downarrow$) 0.0072 0.0050
Table 11: Metrics of 50 generated molecules from GEOM using GeoLDM with neural network guidance (i.e. AIMNet2) on force. * and bold denote the overall best result and our best result, respectively. Percentage changes between our results and GeoLDM are shown in parentheses. For ease of comparison, we report the results using the best scale of ChemGuide from Table 2.
Metric Guidance Scale ChemGuide Baseline
0.0001 0.001 0.01 0.1 1.0 scale = 0.1 EDM GeoLDM
Force RMS (Eh/Bohr) 0.0418 0.0418 0.0418 0.0409 0.0406 (-15.14$\downarrow$) 0.0411 (-14.16$\downarrow$) 0.0742 0.0478
Validity (xTB) 50.00% (1.00$\uparrow$) 50.00% 50.00% 50.00% 50.00% 50.40% (1.40$\uparrow$) 46.40% 49.00%
Uniqueness 100.00% 100.00% 100.00% 100.00% 100.00% 100.00% 100.00% 100.00%
Atom Stability 85.01% 85.01% 85.01% 85.13% 85.26% (0.73$\uparrow$) 84.36% (-0.17$\downarrow$) 81.22% 84.53%
Energy above ground state (Eh) 0.2392 0.2392 0.2392 0.2356 0.2299 (2.26$\uparrow$) 0.1935(-13.92$\downarrow$) 0.3742 0.2248
H.4 Generalization Analysis of Neural Regressors